OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x)=(x-7)/(x-3) and g(x)=(-3x-7)/(x-1)

hartnn (hartnn):

hi :) do you know how to find f(g(x)) ?

OpenStudy (anonymous):

umm I kind of remember so I have to plug in g in f(x)? Or is it the opposite

hartnn (hartnn):

plug in g(x) in f(x)

OpenStudy (anonymous):

Okay so do I solve [(-3x-7)/(x-1)-7]/[((-3x-7)/(x-1)-3]?

hartnn (hartnn):

yes! go ahead :)

OpenStudy (anonymous):

:) I got 5x/3x+2

hartnn (hartnn):

how did you not get that = x ?? show your work...

hartnn (hartnn):

numerator : (-3x-7)-7(x-1) = -3x-7-7x +7 = -3x-7x = -10 x

hartnn (hartnn):

denominator : -3x-7 +3 (x-1) = -3x-7 +3x -3 = -10 actually, the inverse must be different, check your question, there is a sign mistake somewhere.

OpenStudy (anonymous):

Yes! it's x+3!

hartnn (hartnn):

i thought so :)

OpenStudy (anonymous):

Sorry I'm trying to work this out I'm a bit slow

hartnn (hartnn):

sure, take your time, after you finish , verify with these steps, numerator : (-3x-7)-7(x-1) = -3x-7-7x +7 = -3x-7x = -10 x denominator : -3x-7 +3 (x-1) = -3x-7 +3x -3 = -10 num / denom = -10x/ (-10) = x

OpenStudy (anonymous):

For the numerator I got: (-10x-8)/(x-1) ... Do you know what I'm doing wrong?

hartnn (hartnn):

if you show all your steps, i'll help spot the error.

hartnn (hartnn):

-10x is correct, /(x-1) is also correct, only there shouldn't be -8

OpenStudy (anonymous):

I took the -7 and added x-1 to the top and bottom and then put the two fractions together to get: (-3x-7-7x-1)/(x-1) and then I simplified the top

hartnn (hartnn):

-7(x-1) = -7x - 7(-1) = -7x +7

hartnn (hartnn):

so, -3x-7-7x+7 =... ?

OpenStudy (anonymous):

-10x! :)

OpenStudy (anonymous):

Thank now I'll get the denominator!

hartnn (hartnn):

but did u understand how ??

OpenStudy (anonymous):

Yeah! I just accidentally kept the 7 a negative even though I was multiplying two negatives!

OpenStudy (anonymous):

Oh I see no I didn't distribute the 7 to the -1

hartnn (hartnn):

yes! that was your error....

OpenStudy (anonymous):

For the denominator I got -10x/x-1

hartnn (hartnn):

-10 x or just -10 ?

OpenStudy (anonymous):

Oh yeah I meant -10 :) because the -3x+3x=0 so there's no x

hartnn (hartnn):

so now you got f(g(x)) = x ? i think now you can do g(f(x)) =x also :)

OpenStudy (anonymous):

I don't really know what the question means though! lol

OpenStudy (anonymous):

Oh f(g(x)) is just x! And okay so for this question I would just show my work to prove that they both equal x?

hartnn (hartnn):

if \(f(x)=g^{-1}(x), \\then , g(f(x))=g^{-1}(g(x))=x\) the question only asks you to verify this for given pair of functions. and yes, you showed f(g(x))=x already. now just show g(f(x))=x similarly. this would confirm, f and g are inverse of each other,,,,

OpenStudy (anonymous):

Okay I got it now! Thank you for your help and patience! haha :)

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

Sorry @hartnn I'm having trouble getting x! I got -10x/x+3 for the numerator and -10/x+3 for the denominator

hartnn (hartnn):

just divide numerator / denominator x+3 gets cancelled , -10x/ (-10) = x

OpenStudy (anonymous):

Whoops dumb mistake :P thanks again.

OpenStudy (anonymous):

Hey, this is such an old post but it is the same exact question that I have asked, but there is something that bothers me cause I don't quite understand and was hoping you could clarify. numerator : (-3x-7)-7(x-1) = -3x-7-7x +7 = -3x-7x = -10 x denominator : -3x-7 +3 (x-1) = -3x-7 +3x -3 = -10 How come the (x-1) is by the -7 when I thought it was part of the fraction in g(x) shouldn't you divide it in the numerator. I'm only asking this cause you seem to be the only one to help with difficult questions lol and I also just want to learn it so if I happen to come across it later on.

hartnn (hartnn):

I multiplied numerator and denominator by (x-1) so that fractions get eliminated from numerator and denominator. \(\Large \dfrac{(\dfrac{-3x-7}{x-1})-7}{(\dfrac{-3x-7}{x-1})+3} \times \dfrac{x-1}{x-1} \) \(\Large \dfrac{(\dfrac{-3x-7}{\cancel{x-1}}) \times \cancel {(x-1)}-7(x-1)}{(\dfrac{-3x-7}{\cancel{x-1}}) \times \cancel{(x-1)}+3(x-1)} \) \(\Large \dfrac{(-3x-7)-7(x-1)}{(-3x-7)+3(x-1)} \) Hope this helps :)

hartnn (hartnn):

@CrzyWlfMedina ask if you still have any doubt in any step :)

OpenStudy (anonymous):

OMG THANK YOU SOOOO MUCH :D I can't believe I missed something so simple xD. You are awesome @hartnn